OpenStudy (anonymous):

a town planner creates plans for a new water tower. tower is 96 feet tall and the tank is circular of 64 feet 1. the engineer builds a model of the tower of 1 inch : 8 feet. what are the dimensions of the model? 2. supposed the engineer decides to build a second model such that the height of model is 30 inches. what is the scal for the model? @Directrix

Directrix (directrix):

I think there is a missing word here: tank is circular of 64 feet. Tank is circular what of 64 feet.

OpenStudy (anonymous):

diameter

Directrix (directrix):

I'm not understanding: Tank is circular diameter of 64 feet? Is there a diagram or picture of the tank in your book? What shape is the tank? Is is a right circular cylinder? Is it a sphere atop supports? I just need more information.

OpenStudy (anonymous):

the tank is circular WITH a diameter of 64 feet. sorry im just really tired

Directrix (directrix):

No problem with that.

Directrix (directrix):

The tank and the model will be similar figures or solids. I'm thinking the engineer is drawing the model on paper. x = model tower height in inches 1in is to 8 ft as x in is to 96 ft 1/8 = x/96 x = ?

OpenStudy (anonymous):

x=12?

Directrix (directrix):

Yes.

Directrix (directrix):

y = model tank diameter in inches 1/8 = y/64 y = 8 inches We already have x = 12 inches. So, that is part 1.

OpenStudy (anonymous):

so is the answer the model is 12 inches tall and 5 1/3 inches in diameter and the scale of the is 1 inched : 3 1/6 feet?

Directrix (directrix):

Are you saying you think that is the answer for part 2?

OpenStudy (anonymous):

yes

Directrix (directrix):

what is the scale for the model? 30 inches to 96 feet is 5 inches to 16 feet. One inch on the model represents 16/5 feet on the real tower.

Directrix (directrix):

Are your answer options in feet and inches or all in inches or all in feet?

OpenStudy (anonymous):

feet and inches

Directrix (directrix):

Then, one inch on the model is 3.2 feet on the actual tower.

Directrix (directrix):

1. the engineer builds a model of the tower of 1 inch : 8 feet. what are the dimensions of the model? Here, we got 12" for the height of the model. The circular base is 8 inches in diameter? Did we check that?

Directrix (directrix):

2. 2. supposed the engineer decides to build a second model such that the height of model is 30 inches. what is the scal for the model? This ^^ just asks for the scale. Then, one inch on the model is 3.2 feet on the actual tower. Did we agree on that?

OpenStudy (anonymous):

yes

Directrix (directrix):

>>>so is the answer the model is 12 inches tall and 5 1/3 inches in diameter <-- I got 8 inches for the diameter. You have 5 1/3 and (2) the scale of the is 1 inch : 3 1/6 feet? <--- I got 3 1/5 here where you got 3 1/6.

Directrix (directrix):

Do we need to re-work any of this?

Directrix (directrix):

If you post the options on the thread, I can see them. After we match our answers to the correct option, you can delete the options.

OpenStudy (anonymous):

answer choice a is the model is 8 inches tall 6 1/3 inches in diameter. the scales of the second is 1 inch : 3 1/2 feet answer choice b is the model is 12 inches tall and 5 1/3 inches in diameter the scale of the second model is 1 inch : 3 1/6 feet. answer choice c is the model is 14 inches tall and 5 1/2 inches in diameter. the scale of the second model1 inch : 38 feet

Directrix (directrix):

This one is the closest to being correct that I see: answer choice b is the model is 12 inches tall and 5 1/3 inches in diameter the scale of the second model is 1 inch : 3 1/6 feet.

Directrix (directrix):

Would you look and see if there is a typo in the diameter of the tower >>>tank is circular of 64 feet. Is the diameter supposed to be 64 feet?:

OpenStudy (anonymous):

yes

Directrix (directrix):

Is there an option D?

Directrix (directrix):

If not, I would have to go with B if it were my work. I think there's a typo in the printing of the problem.

OpenStudy (anonymous):

no

Directrix (directrix):

The tower model is 12". We can't go against that.

OpenStudy (anonymous):

 are they similar?

Directrix (directrix):

Are there any marked parallel segments?